OpenStudy (tiff9702):

Solve for x . x/3 = 3 1/4

OpenStudy (tiff9702):

@nonopro

OpenStudy (anonymous):

well, 4.x/3=31 4x=31/3 x=31/3.1/4 x=31/12

OpenStudy (mrnood):

@Mrs_Allison I think you have misread the question @tiff9702 is it \[\frac{ x }{ 3 }= 3\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

Oh!!!! im sorry, ill do it again for you

OpenStudy (tiff9702):

@MrNood Yes it is

OpenStudy (mrnood):

multiply both sides by 3 that will give you the answer immediately

OpenStudy (anonymous):

x=9/4

OpenStudy (anonymous):

x/3=3.1/4 4x/3=3 4x=9 x=9/4

OpenStudy (mrnood):

@Mrs.Allison That is not correct

OpenStudy (anonymous):

how so?

OpenStudy (mrnood):

multiply 3 and 1 quarter by 3

OpenStudy (mrnood):

@tiff9702 can you do this? multiply BOTH sides of the equation by 3 @Mrs.Allison Please leave the answer for asker to try

OpenStudy (anonymous):

what i did its correct, im pretty sure. I´m a math teacher --´

OpenStudy (mrnood):

Well - I assure you it is very wrong

OpenStudy (mrnood):

@Mrs.Allison If oyu wan to do it your way (which is more complicated than mine) then first youmust convert the right side to improper fraction 13/4

OpenStudy (anonymous):

\[\frac{ x }{ 3 } = 3* \frac{ 1 }{ 4 }\]\[\frac{ x }{3 }= \frac{ 3 }{ 4 }\]\[x = \frac{ 3 }{ 4 } * 3\]\[x = \frac{ 9 }{ 4 }\]

OpenStudy (anonymous):

did your answer was the same as mine?

OpenStudy (anonymous):

calm yourself down please. And yes, if you multiply one side of the equation by 3 you have to multiply the other one too

OpenStudy (mrnood):

@Mateaus I'm afraid your answer isALSO wrong PLEASE just simply multiply BOTH sides by 3 x/3 *3 = x (3 and 1/4) * 3 =????

OpenStudy (anonymous):

isn't the main expression \[\frac{ x }{ 3 } = 3 * \frac{ 1 }{ 4 }\]??

OpenStudy (mrnood):

no - there is no multiply on the right side it is 3 and 1/4 = (13/4)

OpenStudy (anonymous):

how did you get 13?

OpenStudy (mrnood):

3 = 12/4 + 1/4 = 13/4

OpenStudy (anonymous):

there is no a + which is why I assumed it was multiplication

OpenStudy (mrnood):

but it is the normal way of writing a whole number and fraction

OpenStudy (anonymous):

What? I'm confused, if the 3 is multiplying the 1/4 then my answer is correct, it is not, then I don't know what is that 3 suppose to mean, there's no signal

OpenStudy (anonymous):

Then it is 3 multiplied by 1/4

OpenStudy (anonymous):

@tiff9702 What is the sign between 3 and 1/4?

OpenStudy (mrnood):

NO it's not!! Just SAY it "" three and a quarter" like "two and five eights" or "six and three sixteenths" it is \[3\frac{ 3 }{ 4}\]

OpenStudy (mrnood):

sorry \[3\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

so...

OpenStudy (mrnood):

so \[\frac{ x }{ 3 } = \frac{ 13 }{ 4 }\] so \[x=\frac{ 39 }{ 4} = 9\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

Well according to tiff there is not a sign so we can assume it is how MrNood is describing it.

OpenStudy (anonymous):

There's no sign?

OpenStudy (mrnood):

@tiff9702 I expect you have lost interest by now My answer above is correct as you confirmed my original post of your equation.

OpenStudy (anonymous):

She just told me it looks the same as how MrNood is showing.

OpenStudy (tiff9702):

lol

OpenStudy (anonymous):

Solving problems mhmm hmmm